the larger of two numbers is 1 more than 3 times the smaller. the difference between 8 times the smaller and 2 times the larger is 10. find the numbers.

Answer by checkley71(8403)

let x be the smaller number & y the larger number.
y=3x+1
8x-2y=10
now substitute (3x-1) for y & solve for x
8x-2(3x-1)=10
8x-6x+2=10
2x=10+2
2x=12
x=12/2
x=6 answer for the smaller number.
y=3*6+1
y=18+1
y=19 answer for the larger number.
proof:
8*6-2*19=10
48-38=10
10=10